### v2

# Overcoming data challenges to measure whole-person health in electronic health records

Abstract
Objective: The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning. Data from electronic health records (EHR) present a huge opportunity to operationalize the ALI in the learning health system, except they are prone to missingness and errors. Validation of EHR data (e.g., through chart reviews) can provide better-quality data, but realistically, only a subset of patients’ data can be validated, and most protocols do not recover missing data.
Methods: Using a representative sample of $1000$ patients from the EHR at an extensive learning health system ($100$ of whom could be validated), we propose methods to design, conduct, and analyze statistically efficient and robust studies of the ALI and healthcare utilization. With semiparametric maximum likelihood estimation methods, we robustly incorporate all available patient information (validated or unvalidated) into statistical models. Using targeted design strategies, we examine ways to select the most informative patients for validation. Incorporating clinical expertise, we devise a novel validation protocol to promote the quality and completeness of EHR data.
Results: Validation of the EHR data uncovered relatively low error rates and recovered some missing data through auxiliary information in patients’ charts. Through bespoke simulation studies based on preliminary data, residual sampling was identified as the most informative strategy for completing our validation study. Statistical models incorporating the partial validation data indicated higher odds of engaging in the healthcare system were associated with worse whole-person health (i.e., higher ALI), adjusting for age.
Conclusion: Targeted validation with an enriched protocol allowed us to ensure the quality and promote the completeness of the EHR. Findings from our validation study were incorporated into analyses as we operationalize the ALI as a whole-person health measure intended to predict healthcare utilization in the academic learning health system.
Keywords chart review  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ outcome-dependent sampling  $\cdot$ two-phase designs
1 Introduction
1.1 Measuring Whole-Person Health
Learning healthcare systems are transitioning from a reactive (treatment) to a proactive (prevention) philosophy. This transition could prevent future healthcare utilization through early intervention, improving health for the individual and providing cost savings to the institution. Having a standardized summary measure of health that is easy to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
Numerous data types exist within the electronic health record (EHR), including vital signs, demographics, diagnoses, procedures, and clinical notes. Easy access to a standardized and valid summary measure of whole-person health in the EHR would be advantageous for patients, clinicians, and researchers [1, 2]. The electronic frailty index serves this purpose but is only designed for those aged 65 and older [3]. The availability of a whole-person health metric for the general adult population (age $\geq 18$) would allow for longitudinal tracking of individual and population health within a healthcare system, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
Though individual physiological variables like systolic blood pressure measure aspects of whole-person health, they do not provide a comprehensive assessment across various organ systems. The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [4, 5, 6]. The ALI has been used to predict numerous health outcomes (e.g., [7, 8, 9, 10]) and measure racial disparities in health [11]. Because it is sensitive to short-term interventions, the ALI is also an appealing “surrogate” outcome to test health interventions, like lifestyle and environmental modifications.
Many ALI formulations exist. We focus on the calculation from Seeman et al. (2001) [5], which computes the ALI from ten component measurements across three body systems. As illustrated in Supplemental Figure S1, the original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically driven thresholds to define binary “stressor” indicators (e.g., $=1$ if hemoglobin A1c $\geq 6.5$ and $=0$ if hemoglobin A1c $<6.5$). The ALI is ultimately calculated as the sum of these ten stressors, i.e., the number of “unhealthy” measurements the patient had, with values of zero and ten representing the best and worst whole-person health, respectively.
1.2 Data Quality Challenges in Electronic Health Records Data
EHR data present a huge opportunity to operationalize the ALI in the learning health system. The volume of routinely collected data in the EHR is steadily climbing, increasing the accessibility to clinically meaningful variables. Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains, including HIV and AIDS [12], genetics [13], emergency medicine [14], and therapeutic effectiveness [15].
However, despite clear advantages, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors. These quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [16, 17, 18, 19]. In calculating the ALI, the original measurements may be (i) missing, making the stressor missing, or (i) non-missing but mismeasured, potentially misclassifying the stressor.
Consider a patient whose true underlying HbA1c was $6.5\%$, such that their corresponding stressor should equal one. We illustrate two scenarios for this patient’s data in Figure 1. In the first, they had HbA1c measured, but the value was erroneously transposed to $5.6\%$. Based on the EHR, this patient’s stressor for HbA1c would be misclassified as zero since $5.6\%$ falls into the healthy range. In the second, they did not have HbA1c measured, so the value is missing, and we cannot define a stressor. Simply treating missing components as zero would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining the ALI as the proportion of non-missing stressors a patient experienced.
Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation. In the top scenario, the patient had HbA1c measured but with error, and the correct value was recovered through validation. In the bottom scenario, the patient did not have HbA1c measured, but auxiliary information was located through validation to replace the missing component.
Ignoring errors and missingness in the EHR can hinder analyses, as biased estimates may lead to invalid inference and incorrect conclusions [20, 21]. Errors in EHR data can also reduce the accuracy of prediction models [22]. There is a particular concern for missingness and errors in derived variables, like the ALI, that combine data from many different variables [23, 24]. Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about the association between ALI and healthcare utilization.
For the ALI to be useful in research, it must be reliably and accurately measured. The original calculations, like Seeman et al. (2001) [5], were done in the context of clinical trials, where all patients must have all measurements and data validation protocols are likely in place to ensure quality. Similarly, validation, wherein the extracted EHR data are compared to original source documents (like the electronic chart) provides insights into data quality. Complete data validation (i.e., for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
Fortunately, partially validating EHR data can help identify issues and pave the way for corrections. It provides a cost-effective alternative to complete data validation and is a promising approach to gauge the quality of observational data sources [25, 26, 27]. These partial validation studies are a special case of two-phase designs [28, 29] (Figure 2A). Phase I includes variables available for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation in Phase II.
Figure 2: Diagram of two-phase designs with $N$ patients extracted from the electronic health records (EHR) in phase I and $n$ patients chosen for validation via chart review in phase II. The least informative strategy to select the validation study is A) simple random sampling. More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model. Within strata, the darker shaded areas indicate which patients would be selected.
Validated data are usually expected to be more correct (i.e., less error-prone) but not more complete (i.e., less missingness). To calculate the ALI from EHR data, we needed validation to do both. Herein, we devise a novel validation protocol that incorporates a “roadmap” created by clinical experts to recover missing data created by clinical experts using auxiliary health information (e.g., related diagnoses).
System Component Threshold Auxiliary Information
Cardiovascular Systolic Blood Pressure $>140$ Hypertension
Diastolic Blood Pressure $>90$ Hypertension
Metabolic Body Mass Index $>30$ Obesity; Morbid Obesity;
Grade I, II or III Obesity
Triglycerides $\geq 150$ Hypertriglyceridemia
Total Cholesterol $\geq 200$ Hypercholesterolemia
Inflammation C-Reactive Protein $\geq 10$ Sepsis; Infection; Auto-Immune
Inflammatory Syndrome
Hemoglobin A1C $\geq 6.5$ Diabetes; Impaired
Glycemic Control
Serum Albumin $\geq 3.5$ (None Given)
Creatinine Clearance $<110$ (Males) Renal Failure;
$<100$ (Females) Insufficiency; Acute
Kidney Injury;
Chronic Renal Failure
Homocysteine $>50$ Hyperhomocysteinemia;
Vitamin deficiency
Table 1: Ten component stressors of the allostatic load index were defined by discretizing measurements across three body systems at clinically driven thresholds. If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic. If the auxiliary information was present, the component was treated as a “yes” and included in the patient’s validated ALI.
1.3 Overview
While very insightful, partial validation studies are expensive, time-intensive undertakings. Using a representative sample of $1000$ patients from the EHR at an extensive academic learning health system ($100$ of whom could be validated), we propose novel methods to design, conduct, and analyze statistically efficient and robust studies of whole-person health (ALI) and healthcare utilization. Using advanced statistics and bioinformatics methods, we seize three key opportunities to make the most of our investment: (i) robust and efficient statistical modeling, (ii) enriched validation protocol, and (iii) targeted validation study design.
Statement of Significance
Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in the EHR to make patient- and institution-level decisions. However, routinely collected data are prone to missingness and errors.
What Is Already Known Validation, wherein the extracted EHR data are compared to the original clinical source documents (like the electronic chart), provides insights into data quality, but validating all patients is not feasible, and existing protocols can only correct for errors, not missingness.
What This Paper Adds Building on the statistics and bioinformatics literature, we 1) devise a robust and efficient model for partially validated, error-prone EHR data; 2) propose a targeted validation study design to identify the most informative patients for our model; and 3) develop an enriched validation protocol that incorporates clinical expertise to improve the quality and completeness of EHR data.
2 Methods
2.1 Model and Data
This study was approved by the Institutional Review Board at Wake Forest University School of Medicine. We are interested in modeling healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period. The exposure of interest is the ALI ($X\in[0,1]$), defined as the proportion of stressors a patient experienced. Further, we control for patient age at first encounter ($Z\in\mathcal{R}$).
These variables are assumed to be related through the usual logistic regression model for the log-odds that $Y=1$ (engaged in the healthcare system) given $X$ and $Z$. A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (ORs), denoted ${\boldsymbol{\beta}}$. Due to missingness and errors in the EHR, only an error-prone surrogate measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available. We compute $X^{*}$ as the proportion of non-missing stressors the patient experienced, according to the EHR data. Age $Z$ is fully observed and expected to be error-free.
To obtain consistent estimates for the desired healthcare utilization model, a two-phase design is planned (Figure 2A). In Phase I, the outcome $Y$ and surrogate exposure $X^{*}$ will be collected from the EHR for all $N$ patients. In Phase II, these data will be augmented with the validated exposure $X$ from a chart review for a subset of $n$ patients ($n<N$). To distinguish validated and unvalidated patients, let $V_{i}$ be a validation indicator, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise.
Incorporating data from Phases I and II, along with the validation indicator, the joint distribution of a complete observation is assumed to be
$\Pr(V,X^{*},Y,X,Z)$ $=\Pr(V\mid Y,X^{*},Z){\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)\Pr(X\mid X^{*},Z)\Pr(X^{*},Z),$ (1)
where $\Pr(V\mid Y,X^{*},Z)$ is the validation sampling probability; ${\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)$ corresponds to the logistic regression model; $\Pr(X\mid X^{*},Z)$ is the exposure error mechanism; and $\Pr(X^{*},Z)$ is the joint probability density or mass function of $X^{*}$ and $Z$. Two assumptions are made in defining (1). First, selection for validation (i.e., $V$) is assumed to depend only on Phase I variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [30]. Second, the exposure $X^{*}$ from the EHR is assumed to be a surrogate for the validated exposure $X$, such that $Y$ and $X^{*}$ are conditionally independent given $X$. In other words, once $X$ is known, $X^{*}$ is assumed not to contribute anything further about the relationship with $Y$. While other ways of factoring (1) are possible, this factorization avoids making an additional assumption about $\Pr(X^{*},Z)$.
2.2 Robust Statistical Model for Partially Validated Data
The observed-data log-likelihood function (hereafter, log-likelihood function) for the log ORs ${\boldsymbol{\beta}}$ is defined based on the joint distribution in (1), but first, two simplifications can be made. First, the validation sampling probability $\Pr(V|Y,X^{*},Z)$ can be omitted from the log-likelihood for ${\boldsymbol{\beta}}$ because the Phase II variables are MAR. Second, the joint probability $\Pr(X^{*},Z)$ can be ignored because it will drop out as the log-likelihood function is maximized with respect to ${\boldsymbol{\beta}}$ since $X^{*}$ and $Z$ are fully observed.
Observations for all $N$ patients in the study are assumed to be independent and identically distributed (i.i.d.) following (1). Now, the log-likelihood function for ${\boldsymbol{\beta}}$ is proportional to the following:
$\sum_{i=1}^{N}V_{i}\left[\log\left\{{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid X_{i},Z_{i})\right\}+\log\left\{\Pr(X_{i}\mid X_{i}^{*},Z_{i})\right\}\right]$
$+\sum_{i=1}^{N}(1-V_{i})\log\left\{\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})\right\}.$ (2)
Notice that unvalidated patients contribute their incomplete information to this log-likelihood by summing over the missing validated exposure $X$:
$\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})$ $\propto\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i},x,Z_{i},X_{i}^{*})={\Pr}_{{\boldsymbol{\beta}}}(Y_{i},Z_{i},X_{i}^{*}),$ (3)
remembering that $X$ takes on values $x\in\{0,0.1,\dots,0.9,1\}$ here.
Still, a model for the exposure error mechanism $\Pr(X\mid X^{*},Z)$ is needed to estimate the analysis model parameters since $X$ is only observed for the subset of validated patients. Since the exposure $X$, representing the patient’s ALI, is bounded between zero and one (inclusive), there are not many flexible, parametric choices to model the distribution of $X$ given ($X^{*},Z$). Thus, a nonparametric estimator is preferable, allowing the validated exposure $X$ to take on any values and arbitrarily depend on $X^{*}$ and $Z$.
Lotspeich et al. (2022) [31] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described. The current setting is a special case with only exposure error, and a brief outline is provided in the Supplemental Materials. For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see Lotspeich et al. (2022) [31].
2.3 Targeted Designs to Select Informative Patients for Validation
As a baseline, consider the simple random sampling (SRS) design, wherein $n$ of the $N$ patients are chosen for validation with equal probability. This strategy ignores all available information, making it the most uninformative. For full-likelihood approaches like the SMLE, any information available for all patients can – and should – be used to design the validation study. Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated log OR on ALI), we start at the end. We begin with the planned logistic regression model of $Y$ given $X$ and $Z$, and work backward to determine which patients will be most informative to it. For this reason, the “best” designs vary across different estimators, with model-based estimators, like the SMLEs, offering the most design flexibility. Other estimators that rely on the design–rather than a model–to handle the missingness, like inverse probability weighting, require nonzero sampling probability for all patients, but model-based estimators do not.
We considered stratified and rank-based validation study designs. Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [32, 26, 33]. Ultimately, simulations based on preliminary validation data (see Section 3.3.3) guided our selection of the targeted design used in the application.
Given our binary outcome, the case-control (CC) design is a natural option here, wherein $n/2$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 2B). The balanced case-control (BCC) design further stratifies on covariate information, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 2C). When the covariates are continuous, they must be discretized to form strata. For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average. Then, $n/4$ patients are validated from each of the four $(Y,X_{D}^{*})$ strata. Importantly, we can still fit the model using the original continuous values.
While the CC and BCC designs are intuitive and can offer some efficiency gains over SRS, there are slightly more complicated designs that are guaranteed to be theoretically “optimal,” meaning that they choose the number of patients to validate from each stratum to achieve the lowest possible variance for the log OR on ALI. Many have been proposed to select patients for the collection of “expensive” binary exposures, like biomarker measurement or validation via chart review, for a variety of analytical approaches to modeling binary outcomes [34, 35, 36, 37, 29, 38]. Recently, Lotspeich et al. (2024) [33] proposed a design for optimal validation of secondary use data for maximum likelihood estimation with binary outcome and exposure misclassification (Figure 2D), which could be applied here. However, we would achieve the lowest possible variance for the discretized version $X_{D}^{*}$ rather than $X^{*}$. Increasing the number of unique values of $X_{D}^{*}$ (i.e., defining more strata) can help, but information will still be lost.
Still, designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains. For example, extreme tail sampling (ETS) designs focus on the extremes (i.e., tails) of the distribution, ranking the data by some value of interest and targeting patients with the smallest or largest values. To identify patients who are most informative not just about ALI ($X^{*}$ and $X$) but about its relationship to healthcare utilization ($Y$), residual sampling (RS) is preferred [39]. With measurement error, we begin by fitting the “naive” analysis model for $Y$ (i.e., using $X^{*}$ instead of $X$) to obtain estimates for the log ORs $\hat{\boldsymbol{\beta}}^{*}$. Then, residuals for the naive model are computed for each patient $i$ ($i=1,\dots,N$) as
$r_{i}^{*}$ $=Y_{i}-{\Pr}_{\hat{\boldsymbol{\beta}}^{*}}(Y_{i}=1|X^{*}_{i},Z_{i})$
$=Y_{i}-\frac{1}{1+\exp\left\{-\left(\beta_{0}^{*}+\beta_{1}^{*}X_{i}^{*}+{\boldsymbol{\beta}}_{2}^{*}Z_{i}\right)\right\}},$ (4)
and used to rank the $N$ patients in the study. With logistic regression, the largest residuals $r_{i}^{*}$ (in magnitude) correspond to observations where the observed value $Y_{i}$ had a low predicted probability (Figure 2E). The auditDesignR package, available on GitHub at www.github.com/sarahlotspeich/auditDesignR, contains R functions to sample from all of these designs.
2.4 Enriched Validation Protocol to Reduce Errors and Missingness
The chart reviews to validate the EHR data were conducted over $6$ months (February to July 2024) by a team of four clinical research assistants and technicians in the Wake Forest University School of Medicine (hereafter, “auditors”). These skilled auditors were chosen for their familiarity with Epic, the institution’s electronic charting software, and their demonstrated experience in the clinical research environment.
2.4.1 Identifying Sources of Auxiliary Information
Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1). While it is unlikely that we would identify the original missing measurements, we harnessed the clinical expertise of our team to identify sources of auxiliary information about the missing discretized components. The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values.
For example, if a patient were missing HbA1c, the auditors would search their chart for a diagnosis of diabetes or impaired glycemic control. If either were present, the patient’s diagnosis would suggest their missing HbA1c value would have exceeded the clinically meaningful threshold of $6.5\%$. Thus, their validated ALI $X$ would include one point for HbA1C. If both were absent, the patient’s validated ALI would continue to exclude this still-missing component.
Many components, like C-reactive protein and HbA1c, had multiple sources of auxiliary information, while others, like triglycerides and total cholesterol, had just one. Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
2.4.2 Validated Data Collection
The EHR data were extracted by the Wake Forest Clinical and Translational Science Institute and securely received by the analytical team. In preparation for validation, the analytical team transformed the data from the original wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable). This transformation streamlined the data structure for the auditors, with only one value to validate per row (Figure 3).
Figure 3: Illustrative example of how the original extracted EHR data for patients’ vitals were transformed from a wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable), merged with the audit roadmap, and augmented with empty columns for auditors to enter their findings. The same was done with patients’ lab data.
Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one found in Epic). Depending on the scenario, the auditors would enter different information for the reviewed value, and each data point was assigned an audit finding (flow chart in Figure 4). Three findings were possible for non-missing components in the EHR.
  1. 1.
    Extracted value correct: If the component matched Epic, the auditor would enter the original value again.
  2. 2.
    Extracted value incorrect: If the component did not match Epic, the auditor would enter the new value found in Epic.
  3. 3.
    Extracted value not found: If the component could not be found in Epic, the auditor would enter “Not Found.”
Two findings were possible for missing components.
  1. 4.
    Auxiliary information found: If auxiliary information from the audit roadmap were found in Epic, the auditor would enter the auxiliary information.
  2. 5.
    Auxiliary information not found: If auxiliary information from the audit roadmap were not found in Epic, the auditor would enter “Not Found.”
Figure 4: Flow chart of instructions given to auditors for the chart reviews (Steps 1–3) and the corresponding audit finding assigned by the analysts (Step 4). Column names refer to those in Figure 3.
3 Results
3.1 Cohort Description
We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina. To be considered, patients needed to (i) be $18-65$ years old and (ii) initiate their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic). Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
Our data contained demographics, vitals, lab values, discharge diagnoses, and problem list values dating $3$ years before the first encounter. Key variables included the measurements for the ten components of the ALI (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period. Since most patients had multiple encounters, the ALI components were based on mean measurements.
Sampled patients had a median age of $48$ years and a median income of \$52,320. They were $61\%$ female, $72\%$ White, $6\%$ Hispanic, and $73\%$ homeowners. There were $318$ patients ($32\%$) with at least one hospitalization or ED visit. Based on the original unvalidated EHR data, the distribution of error-prone ALI was right-skewed (Supplemental Figure S2). With a median of $0.33$ (interquartile range $=0.17,0.50$), on average, patients in our study experienced $1$ out of every $3$ non-missing stressors on their whole-person health.
3.2 Baseline Data Quality and Naive Model Estimates
As expected, the extracted EHR data had a large amount of missingness. In particular, rare labs for inflammation components were missing for $>95\%$ of patients. Labs for the metabolic components were missing for $21\%$ of patients, and measurements for the cardiovascular components were non-missing for basically everyone (Supplemental Figure S3A). The most common patterns of missing data are displayed in Supplemental Figure S4. Fortunately, most patients ($78\%$) had at least $5$ non-missing ALI components (Supplemental Figure S3B).
Using these ALI values to fit the “naive” healthcare utilization model, we estimate the following. First, for every $0.1$-point (or one component) increase in ALI, the odds of engaging in the healthcare system are expected to increase by $10\%$ (OR $=1.10$, 95% confidence interval [95% CI]: $1.03$, $1.18$), adjusting for age. Second, the expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 are $0.25$ (95% CI: $0.18$, $0.35$). Third, for every $10$ years older a patient was at their first encounter, their odds of engaging in the healthcare system are expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI. Thus, the naive model suggests that worse whole-person health leads to higher expected odds of hospitalization or emergency department visits.
3.3 Three Waves of Data Validation Via Chart Review
While the naive model aligns with the existing literature, we wanted to ensure that the missingness and potential errors in the EHR data did not lead to spurious conclusions. Thus, we secured funding to validate data for $n=100$ of the $N=1000$ patients, whose selection and review were broken down into three waves: (i) Pilot, (ii) Wave I, and (iii) Wave II.
3.3.1 Pilot Wave: Testing the Audit Protocol and Establishing Inter-Auditor Agreement
The first $n_{0}=4$ patients were selected for chart review via the BCC design introduced in Section 2.3. Specifically, patients were split into four strata based on their outcome and error-prone ALI, discretized at the median:
  • (i)
    Engaged in the healthcare system and low error-prone ALI ($Y=1$, $X^{*}\leq 0.33$),
  • (ii)
    Engaged in the healthcare system and high error-prone ALI ($Y=1$, $X^{*}>0.33$),
  • (iii)
    Did not engage in the healthcare system and low error-prone ALI ($Y=0$, $X^{*}\leq 0.33$), and
  • (iv)
    Did not engage in the healthcare system and high error-prone ALI ($Y=0$, $X^{*}>0.33$).
Then, $1$ patient was randomly chosen for validation from each stratum.
The auditors followed the protocol in Section 2.4 to (i) verify the non-missing values in the EHR data or note any discrepancies and (ii) locate auxiliary information for missing values. Following the completion of the Pilot wave, the auditors and analysts met to discuss any modifications that would be helpful before proceeding. For example, the third audit finding of “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded. Adding an optional “Notes” column, where auditors could include any additional information they uncovered, also resulted from this discussion. The Pilot and Wave I findings were combined to interpret data quality (see Section 3.3.2).
We started with two auditors on our team. Each reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample. Later, two new auditors joined the team, each reviewing these four patients’ records as well. Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [40], as detailed in the Supplemental Materials. We found high agreement between the auditors’ findings for lab ($\kappa=0.90$) and vitals ($\kappa=0.81$) variables.
3.3.2 Wave I: Collecting Preliminary Audit Data
The next $n_{1}=48$ patients validated were also selected via the BCC design from Section 3.3.1. Since we did not have any prior information about the quality of the EHR data, we sampled more conservatively here to collect preliminary information for a more extreme, targeted design in Wave II. Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [32, 26]. Even if prior information is available (e.g., from a previous audit), validating in multiple waves allows us to relearn how our data may have changed before using all our resources to validate them [33].
[…]
Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined. These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable.
Fewer vitals than labs were found to match Epic ($59\%$ versus $81\%$), with $31\%$ not found. Through validation, we discovered that the EHR data extraction algorithm expected every patient encounter to have vitals. For encounters where vitals were not taken (e.g., physical therapy or optometry visits), values were imputed from other encounters; these imputed values could not be found in Epic by the auditors. Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
Figure 6: Comparison of the discretized allostatic load index components from the unvalidated EHR data to A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
Incorporating the audit roadmap into the validation protocol improved the completeness of the data, as the number of missing values decreased for nearly all components of the ALI (Supplemental Figure S3A). Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of the $52$ validated patients ($44\%$). The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average. Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).
Incorporating the Pilot and Wave I, we re-fit the healthcare utilization model using the SMLE with the first $52$ validated and $948$ unvalidated patients’ data (Table 2). The association between whole-person health and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase in ALI (OR $=1.21$, 95% CI: $1.17,1.25$), adjusting for age. For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$). The adjusted OR for age was relatively unchanged (OR $=1.10$, 95% CI: $1.01,1.21$). Thus, the preliminary SMLE model reaffirms that worse whole-person health is associated with higher expected odds of engaging in the healthcare system.
Odds/Odds
Coefficient Ratio Estimate 95% CI
Naive Analysis (Unvalidated Data)
Intercept $0.251$ $(0.176,0.357)$
ALI $1.099$ $(1.028,1.175)$
Age $1.108$ $(1.000,1.228)$
Preliminary SMLE (Pilot + Wave I Validated Data)
Intercept $0.173$ $(0.115,0.262)$
ALI $1.207$ $(1.167,1.247)$
Age $1.104$ $(1.007,1.211)$
Final SMLE (All Waves Validated Data)
Intercept $0.225$ $(0.156,0.327)$
ALI $1.120$ $(1.088,1.154)$
Age $1.105$ $(1.018,1.200)$
Table 2: Estimates and 95% confidence intervals for coefficients from the naive analysis (logistic regression, no validated data), preliminary sieve maximum likelihood estimator (SMLE, validated data on $52$ patients), and final SMLE (validated data on $100$ patients) models.
3.3.3 Simulations: Using Preliminary Audit Findings to Design Wave II
Before validating the remaining patients, we designed bespoke simulation studies based on the preliminary audit data. These simulations allowed us to explore how potential designs compared under different scenarios for the true underlying data quality. In particular, the design offering the best statistical efficiency for the OR on ALI was desired. Focusing here on the takeaways and our decision-making, details like the data-generating mechanisms can be found in the Supplemental Materials. The five designs described in Section 2.3 were considered: SRS, CC, BCC, OPT, and RS.
Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency (i.e., smaller variance) than the uninformative SRS strategy. The RS design provided the best efficiency under extra low to moderate error settings; the preliminary validation data indicated that we had a perfect TPR $=1.00$ and negligible FPR $<0.01$.
The SMLE $\hat{\beta}_{1}$ was empirically unbiased under full/high data recovery rates ($90-100\%$ missing data recovered through validation) and different validation study designs. While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on the unvalidated EHR data). The RS design offered the highest efficiency, regardless of how much missing data was recovered. Thus, based on extensive simulations considering different error rates and data recovery scenarios, we concluded that the RS design was the most informative way to select the remaining $n_{2}=48$ patients for validation in Wave II.
3.3.4 Wave II: Implementing the Targeted Validation Study Design
For the RS design, each unvalidated patient’s residual was calculated based on the naive model estimates, following (4), as
$r_{i}^{*}=Y_{i}-\frac{1}{1+\exp\left\{-\left(-1.383+0.945X_{i}^{*}+0.103Z_{i}\right)\right\}}.$
Then, patients were ranked by $r_{i}^{*}$, and the $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation. Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero. That is, patients whose observed $Y_{i}$ aligned well with the naive model’s prediction were not considered to be very informative and thus were deprioritized for validation.
A total of $3867$ data points were validated in Wave II ($718$ lab and $3149$ vitals variables), which was fewer than in the Pilot and Wave I. Given that similar numbers of patients were validated, this difference was explained by fewer patient encounters, with between $16$ and $453$ validated data points per patient (median $=65.5$). Overall, the estimated error rates for non-missing components (TPR $=99\%$, FPR $=3\%$) and recovery rate for missing data ($17\%$) in Wave II closely resembled the preliminary ones (Figure 6B). Still, some auditors’ findings differed (Figure 5B). For example, more vitals were found to match Epic ($65\%$). Similar amounts of missing data were recovered using the audit roadmap, spread across all lab variables. Additional findings are in Supplemental Table S1.
3.4 Post-Validation Model of Healthcare Utilization After Validation
Combined findings from all waves indicated that the non-missing EHR data were of excellent quality (TPR $=99.5\%$, FPR $=2\%$), and auxiliary information could be located for $21\%$ of missing components (Figures 5C and 6C). Detailed findings are in Supplemental Table S2. We pooled validated data for the $n=100$ total validated patients with the unvalidated on all $N=1000$ patients from the original EHR data to estimate the SMLE once more (Table 2).
All final model estimates fell between the naive analysis and preliminary SMLEs. For each $0.1$-point increase in ALI, a $12\%$ increase in odds of engaging in the healthcare system was expected (OR $=1.12$, 95% CI: $1.09,1.15$), adjusting for age. The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 00 was $0.23$ (95% CI: $0.16$, $0.33$). Again, older patients were expected to be more likely to engage in the healthcare system, controlling for ALI (OR $=1.11$, 95% CI: $1.02,1.20$). Additional technical details, including R code, are available in the Supplemental Materials.
Ultimately, all three models agreed that worse whole-person health (i.e., higher ALI) was associated with a higher likelihood of being hospitalized or visiting the emergency department, adjusting for age. Our chart reviews uncovered very high quality of the non-missing EHR data and located auxiliary information for some missing data. Without validation, knowing whether we can trust the EHR data would not have been possible. Thus, while the validated and unvalidated models led to the same conclusions here, validating EHR data before using them in research is critical.
4 Discussion
Predicting healthcare utilization is a priority in the learning health system, with important implications at both the patient- and institution-levels. Establishing standardized, valid summary measures of whole-person health, like the ALI, associated with utilization and can scale up to the entire institution is an integral step. While EHR data contain variables to calculate these measures, missing data is a key challenge, and non-missing data are still prone to errors.
Data validation or auditing, often through chart reviews, can gauge data quality, but this time-intensive process is usually only possible for a small subset of patients’ data. Also, existing validation protocols do not overcome missingness. Herein, we propose novel methods to maximize information gained from a partial validation study of $100$ patients from an EHR dataset of $1000$ through the statistical model, validation protocol, and targeted strategy to select patients.
The proposed analytical and design strategies were devised to ensure robust and efficient estimates. Specifically, the statistical model avoids making restrictive assumptions about the error mechanism, relating the unvalidated and validated ALI, while incorporating all available information to achieve precise estimates (i.e., low variance). Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients most informative for the clinical question. Our enriched validation protocol builds on those in the literature (e.g., [41]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
While our roadmap helped recover some missing ALI components, the clinical guidelines were only defined for missing values that were presumed to be a “yes.” If auditors were given additional guidance on what auxiliary information might mean that a missing value is a “no,” the missingness could be further reduced. Others have described missingness in the EHR as selection bias and sought to correct for it using approaches like inverse probability weighting [42]. Applying such approaches to our post-validation data may reduce the impact of residual missingness.
Further, EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [43, 44, 45]. For example, patients seeking care more often may have more diagnoses. Our decision to adapt the Seeman et al. (2001) [5] definition of the ALI to be the proportion of non-missing components could be considered a summary measure. In a recent review, Sisk et al. (2021) [46] found summary measures to be the most common approach to handling informative presence bias. We also used the enriched chart reviews to better understand the missingness.
There are several interesting directions for future work. First, patients could have multiple labs and vitals during the study period. The averages of these values were used to calculate the ALI, and a longitudinal approach to modeling healthcare utilization wherein patients’ ALI can vary over time might provide more information. Second, auditors recorded their findings in a Microsoft Excel workbook. An electronic REDCap [47] interface could be built following published guidelines [48] for ease of use and accuracy. Third, one could use other machine learning techniques (e.g., random forest classification) and explore targeted validation strategies to improve predictive ability (e.g., accuracy) rather than statistical efficiency [49, 50].
5 Conclusion
EHR data provide an exciting opportunity to implement key measures, like whole-person health scores, in the learning health system. However, errors and missingness in these routinely collected data pose challenges in verifying and operationalizing such measures. Validation of even a subset of patients’ data is time- and cost-intensive but provides crucial insights into data quality. Statistical and biomedical informatics methods can maximize information gained from such partial validation.
Acknowledgments
The authors gratefully acknowledge Wake Forest University and Wake Forest University School of Medicine for an intercampus collaborative grant that supported this work. Computations were performed using the Wake Forest University (WFU) High Performance Computing Facility, a centrally managed computational resource available to WFU researchers including faculty, staff, students, and collaborators.
Supplementary Materials