### v4

# Overcoming data challenges through enriched validation and targeted sampling to measure whole-person health in electronic health records

Abstract
Objective: The allostatic load index (ALI) is a 10-component composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning. Data from electronic health records (EHR) present a huge opportunity to operationalize the ALI in learning health systems; however, these data are prone to missingness and errors. Validation (e.g., through chart reviews) can provide better-quality data, but realistically, only a subset of patients’ data can be validated, and most protocols do not recover missing data.
Methods: Using a representative sample of $1000$ patients from the EHR at an extensive learning health system ($100$ of whom could be validated), we propose methods to design, conduct, and analyze statistically efficient and robust studies of ALI and healthcare utilization. Employing semiparametric maximum likelihood estimation, we robustly incorporate all available patient information into statistical models. Using targeted design strategies, we examine ways to select the most informative patients for validation. Incorporating clinical expertise, we devise a novel validation protocol to promote EHR data quality and completeness.
Results: Chart reviews uncovered few errors ($99\%$ matched source documents) and recovered some missing data through auxiliary information in patients’ charts. On average, validation increased the number of non-missing ALI components per patient from $6$ to $7$. Through simulations based on preliminary data, residual sampling was identified as the most informative strategy for completing our validation study. Incorporating validation data, statistical models indicated that worse whole-person health (higher ALI) was associated with higher odds of engaging in the healthcare system, adjusting for age.
Conclusion: Targeted validation with an enriched protocol enabled us to ensure the quality and promote the completeness of the EHR. Findings from our validation study were incorporated into analyses as we operationalize the ALI as a scalable whole-person health measure predictive of healthcare utilization in the learning health system.
Keywords chart review  $\cdot$ computable phenotype  $\cdot$ data audits  $\cdot$ healthcare utilization  $\cdot$ measurement error  $\cdot$ missing data  $\cdot$ residual sampling  $\cdot$ two-phase designs
1 Introduction
1.1 Measuring Whole-Person Health
The advent of learning health systems [1, 2] signals a transition from a reactive (treatment-based) to a proactive (prevention-based) philosophy in healthcare. This transition could prevent future healthcare utilization through early intervention, improving health for the individual and providing cost savings to the institution. Having a standardized summary measure of health that is simple to compute and track could play a central role in disease prevention (saving lives) and reducing healthcare utilization (saving costs).
Numerous data types exist within the electronic health record (EHR), including vital signs, demographics, diagnoses, procedures, and clinical notes. Easy access to a standardized and valid summary measure of whole-person health in the EHR would be advantageous for patients, clinicians, and researchers [3, 4]. The electronic frailty index (EFI) serves this purpose but is only designed for those aged 65 and older [5]. A whole-person health metric for adults aged $18+$ would enable system-wide longitudinal tracking of individual and population health, comparisons across geographies and hospitals, and integration of an important predictor in algorithms relevant to the enterprise (e.g., predicting readmissions).
Individual physiological variables measure aspects of health; they do not provide a comprehensive assessment across various organ systems. The allostatic load index (ALI) is a composite measure of whole-person health, which reflects the multiple interrelated physiological regulatory systems that underlie healthy functioning [6, 7, 8, 9]. It has been used to predict numerous health outcomes (e.g., [10, 11, 12, 13]) and measure racial disparities in health [14]. Because it is sensitive to short-term interventions, ALI is also an appealing “surrogate” outcome to test health interventions.
Many ALI formulations exist. We focus on the calculation from [7], which computes ALI from ten components across three body systems (Table 1). The original numeric measurements (e.g., hemoglobin A1c [HbA1c]) are discretized at clinically-driven thresholds to define binary indicators of being at an unhealthy level (e.g., $=1$ if HbA1c $\geq 6.5$ and $=0$ if HbA1c $<6.5$). ALI is ultimately calculated as the sum of these ten components, with zero and ten representing the best and worst whole-person health, respectively (Supplemental Figure S1).
Whole-person health is a latent concept influenced by genetics, environment, and other factors. Scores like the ALI incorporate components across body systems and yield “computable phenotypes” for it, concisely characterizing patients’ current health status with one-number summaries of observed data [15, 16]. The ALI’s ability to quantify whole-person health has been rigorously examined in controlled studies [6, 7, 8]. However, before adopting the ALI in other settings, it must be revalidated, accounting for real-world data challenges.
1.2 Data Quality Challenges in Electronic Health Records Data
EHR data present a substantial opportunity to operationalize the ALI in the learning health system. The volume of routinely collected data in the EHR is steadily climbing, increasing the accessibility to clinically meaningful variables. Analyzing data that come at little to no additional cost for collection is particularly appealing to biomedical researchers, leading to a surge in the utilization of EHR data in many clinical domains (e.g., [17, 18, 19, 20]).
However, despite clear advantages, EHR data are collected in the fast-paced world of clinical care, leading to missingness and errors. Quality concerns remain a significant hurdle to the large-scale adoption of EHR in healthcare research [21, 22, 23, 24]. In calculating the ALI, measurements may be (i) missing, making the component missing, or (ii) non-missing but mismeasured, potentially misclassifying the component.
Consider a patient whose true HbA1c was $6.5\%$, so their corresponding component should equal one. We illustrate two scenarios for this patient’s data in Figure 1. In the first, HbA1c was measured but erroneously transposed to $5.6\%$. Based on the EHR, their HbA1c component would be misclassified as zero since $5.6\%$ falls into the healthy range. In the second, HbA1c was not measured, so the value is missing, and we cannot define a component. Counting missing components as zeros would also induce misclassification by treating them as “healthy." We accommodate missing data by instead defining ALI as the proportion of non-missing components a patient experienced.
Figure 1: Diagram of potential errors and missingness in a patient’s hemoglobin A1c (HbA1c) measurement in the electronic health record (EHR) before and after validation (chart review). In the top scenario, the patient had HbA1c measured but with error, and the correct value was recovered through validation. In the bottom scenario, the patient did not have HbA1c measured, but auxiliary information was located through validation to replace the missing component.
Ignoring errors and missingness in the EHR can hinder analyses, as biased estimates may lead to invalid inference and incorrect conclusions [25, 26]. Errors in EHR data can also reduce the accuracy of prediction models [27]. There is a particular concern for missingness and errors in derived variables, like ALI, that combine data from many different variables [28, 29]. Analyzing the ALI computed from the EHR data “as is” may misrepresent a patient’s actual health and lead to false conclusions about its association with healthcare utilization.
Applying popular missing data approaches like multiple imputation (MI) may offer some improvement at baseline [30]. However, they can also hinder results if any imputation models are misspecified [31]. Since components are likely missing not at random in EHR data (e.g., clinicians usually order labs when they suspect values are abnormal) [32], correctly specifying imputation models could be especially challenging. Plus, MI would not correct for errors in non-missing ALI components.
For the ALI to be useful in patient care and research, it must be reliably and accurately measured. The original calculations, like [7], were done in clinical trials, where all patients must have all measurements, and data validation protocols are likely in place to ensure quality. Similarly, validation via chart review, comparing extracted EHR data to source documents (like electronic charts), provides data quality insights. Complete data validation (for all patients and variables) would ensure the integrity of EHR data but is overly expensive and unattainable in practice.
Fortunately, partially validating EHR data can help identify issues and pave the way for corrections. It provides a cost-effective alternative to complete data validation and is a promising approach to gauge the quality of observational data sources [33, 34, 35]. These partial validation studies are a special case of two-phase designs [36, 37]. Phase I includes EHR data for all $N$ patients, which can be used to select a subsample of $n$ patients ($n<N$) for validation (chart reviews) in Phase II.
Validated data are usually expected to be more correct (i.e., less error-prone) but not more complete (i.e., less missingness). To calculate the ALI from EHR data, we needed validation to do both. We devise a novel validation protocol that incorporates a “roadmap” created by clinical experts to recover missing data using auxiliary information (e.g., related diagnoses). With it, we verify the ALI’s strength as a computable phenotype for whole-person health and a predictor of healthcare utilization in observational settings with data challenges.
System Component Threshold Auxiliary Information
Cardiovascular Systolic Blood Pressure $>140$ Hypertension
Diastolic Blood Pressure $>90$ Hypertension
Metabolic Body Mass Index $>30$ Obesity; Morbid Obesity;
Grade I, II or III Obesity
Triglycerides $\geq 150$ Hypertriglyceridemia
Total Cholesterol $\geq 200$ Hypercholesterolemia
Inflammation C-Reactive Protein $\geq 10$ Sepsis; Infection; Auto-Immune
Inflammatory Syndrome
Hemoglobin A1C $\geq 6.5$ Diabetes; Impaired
Glycemic Control
Serum Albumin $\geq 3.5$ (None Given)
Creatinine Clearance $<110$ (Males) Renal Failure;
$<100$ (Females) Insufficiency; Acute
Kidney Injury;
Chronic Renal Failure
Homocysteine $>50$ Hyperhomocysteinemia;
Vitamin deficiency
Table 1: Ten components of the allostatic load index were defined by discretizing measurements across three body systems at clinically-driven thresholds. If the component was missing from the extracted EHR data, auditors searched for auxiliary information in the patient’s medical chart in Epic (the institution’s electronic charting software). If the auxiliary information was present, the patient was treated as having an “unhealthy” measurement and the component included in their validated ALI.
1.3 Application to the Learning Health System
We investigate how to overcome data challenges to measure whole-person health across one of the largest learning health systems in the United States. Having recently integrated the EFI into the EHR to measure whole-person health for patients $65$ and older [38], the health system is interested in operationalizing the ALI for the broader patient population. We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina. Our sample was drawn from patients (i) $18$-$65$ years old who (ii) initiated their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic). Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
Our data contained demographics, vitals, lab values, discharge diagnoses, and problem lists dating $3$ years before the first encounter. Key variables included measurements for the ten ALI components (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period. Since most patients had multiple encounters, ALI components were based on mean measurements.
1.4 Overview
While very insightful in the development and use of computable phenotypes, partial validation studies are expensive, time-intensive undertakings. Using a representative sample of $1000$ patients from the EHR at an extensive academic learning health system ($100$ of whom could be validated), we propose novel methods to design, conduct, and analyze statistically efficient and robust studies of whole-person health (ALI) and healthcare utilization. Using advanced statistics and bioinformatics methods, we seize three key opportunities to maximize our investment into partial validation: robust and efficient statistical modeling, enriched chart review protocol, and targeted sampling design.
Statement of Significance
  • Problem Standardized and valid summary measures of current health, i.e., whole-person health, are needed in electronic health records (EHR) to make patient- and institution-level decisions. However, routinely collected data are prone to missingness and errors.
  • What Is Already Known Validation, wherein the extracted EHR data are compared to the original clinical source documents (like the electronic chart), provides insights into data quality, but validating all patients is not feasible, and existing protocols can only correct for errors, not missingness.
  • What This Paper Adds Building on the statistics and bioinformatics literature, we 1) devise a robust and efficient model for partially validated, error-prone EHR data; 2) propose a targeted validation study design to identify the most informative patients for our model; and 3) develop an enriched validation protocol that incorporates clinical expertise to improve the quality and completeness of EHR data.
  • Who Would Benefit From the Knowledge in This Paper Biomedical researchers interested in analyzing routinely collected data, like those in the EHR, while correcting their data quality issues.
2 Methods
2.1 Healthcare Utilization Model
This study was approved by the Institutional Review Board at Wake Forest University School of Medicine (WFUSM). We model healthcare utilization ($Y\in\{0,1\}$), defined as a binary indicator of having at least one hospitalization or emergency department (ED) visit in the $2$-year study period. Whole-person health, measured by ALI ($X\in[0,1]$), is the exposure of interest. Each component $S_{j}$ ($j\in\{1,\dots,10\}$) equals one if the patient’s measurement was considered unhealthy and zero otherwise, such that $X=\sum_{j=1}^{10}S_{j}/10$ is the proportion of unhealthy measurements experienced. Further, we control for age ($Z\in\mathcal{R}$).
We assume a logistic regression model for the log-odds that $Y=1$ (patient engaged in healthcare system) given ($X$, $Z$). A sample of $N$ patients was extracted from the EHR to estimate these log odds ratios (logORs), denoted ${\boldsymbol{\beta}}$. Age $Z$ is fully observed and expected to be error-free. However, due to missingness and errors in ALI components, only an error-prone measure for $X$, denoted by $X^{*}$ ($X^{*}\in[0,1]$), is available from the EHR data. We define error-prone components $S_{j}^{*}$ ($j\in\{1,\dots,10\}$) and introduce missingness indicators $M_{j}^{*}$ for them (i.e., $M_{j}^{*}=1$ if $S_{j}^{*}$ is missing). We compute $X^{*}=\sum_{j=1}^{10}(1-M_{j}^{*})S_{j}^{*}/\sum_{j=1}^{10}(1-M_{j}^{*})$, i.e., as the proportion of non-missing components experienced. Calculating ALI as a proportion – rather than sum – was critical. Summing over non-missing components treats missing values as zeros (i.e., healthy measurements); instead, the proportion ignores them.
2.2 Three Key Steps to Partial Validation Studies: Analysis, Design, and Protocol
To obtain consistent estimates for the healthcare utilization model, we planned a partial validation study. We first extracted EHR data ($Y$, $X^{*}$, $Z$) for a sample of $N=1000$ patients. Then, a subset of $n=100$ patients was selected for validation via chart review such that $X$ was also measured.
Partial validation studies break down into three key steps. Design: Select which patients should be validated. Protocol: Develop standardized tools and practices for those patients’ chart reviews. Analysis: Estimate statistical models or quantities of interest from validated and unvalidated data together. These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information. For added flexibility, a multi-wave validation framework, conducting chart reviews in multiple sprints, allows us to revisit key steps and make modifications as information accumulates (Figure 2).
In our application (Section 3), we built the statistical model and developed the baseline sampling strategy and tools before beginning validation. The proposed methods are detailed in order of development: analysis (Section 2.3), design (Section 2.4), and protocol (Section 2.5). Across three validation waves, we updated our sampling strategy and tools (revisiting the design and protocol steps, respectively), before fitting the final model (Section 2.6).
Figure 2: After extracting the electronic health records (EHR) data, partial validation studies break down into three key steps: analysis, design, and protocol. These steps are intertwined, with design choices depending on the planned analysis, and each step provides opportunities to gain information. For added flexibility, a multi-wave validation framework allows us to revisit these key steps and make modifications as information accumulates.
2.3 Analysis: Robust and Efficient Statistical Model for Partially Validated Data
Let $V_{i}$ distinguish validated and unvalidated patients, where $V_{i}=1$ if the $i$th patient is validated, $i\in\{1,\dots,N\}$, and $V_{i}=0$ otherwise. Incorporating EHR and chart review data, along with the validation indicator, the joint distribution of a complete observation is assumed to be
$\Pr(V,X^{*},Y,X,Z)$ $=\Pr(V\mid Y,X^{*},Z){\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)\Pr(X\mid X^{*},Z)\Pr(X^{*},Z),$ (1)
where $\Pr(V\mid Y,X^{*},Z)$ is the validation sampling probability; ${\Pr}_{{\boldsymbol{\beta}}}(Y\mid X,Z)$ is the logistic regression model; $\Pr(X\mid X^{*},Z)$ is the exposure error mechanism; and $\Pr(X^{*},Z)$ is the joint probability density or mass function of ($X^{*}$, $Z$). Two assumptions are made in (1). First, selection for validation depends only on fully observed variables ($Y$, $X^{*}$, $Z$), so the validated exposure $X$ is missing at random (MAR) for unvalidated patients [39]. Second, $X^{*}$ from the EHR is a surrogate for validated $X$ from the chart reviews, such that $Y$ and $X^{*}$ are conditionally independent given $X$.
Observations for all $N=1000$ patients are assumed to be independent and identically distributed following (1). Thus, the observed-data log-likelihood function (hereafter, log-likelihood) for the logORs ${\boldsymbol{\beta}}$ is defined from (1), after two simplifications. First, we can omit $\Pr(V|Y,X^{*},Z)$ from the log-likelihood because $X$ is MAR. Second, we can ignore $\Pr(X^{*},Z)$ because it drops out as the log-likelihood is maximized with respect to ${\boldsymbol{\beta}}$ since ($X^{*}$, $Z$) are fully observed. Now, the log-likelihood for ${\boldsymbol{\beta}}$ is proportional to
$\sum_{i=1}^{N}V_{i}\left[\log\left\{{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid X_{i},Z_{i})\right\}+\log\left\{\Pr(X_{i}\mid X_{i}^{*},Z_{i})\right\}\right]$
$+\sum_{i=1}^{N}(1-V_{i})\log\left\{\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})\right\}.$ (2)
Notice that unvalidated patients contribute incomplete information to (2) by summing over the missing validated $X$:
$\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i}\mid x,Z_{i})\Pr(x\mid X_{i}^{*},Z_{i})$ $\propto\sum_{x=0}^{1}{\Pr}_{{\boldsymbol{\beta}}}(Y_{i},x,Z_{i},X_{i}^{*})={\Pr}_{{\boldsymbol{\beta}}}(Y_{i},Z_{i},X_{i}^{*}),$ (3)
remembering that $X$ takes on values $x\in\{0,0.1,\dots,0.9,1\}$ here.
A model for $\Pr(X\mid X^{*},Z)$ is needed to estimate ${\boldsymbol{\beta}}$ from (2) since $X$ is only observed for validated patients. Since $X\in[0,1]$, there are not many flexible, parametric choices to model its distribution given ($X^{*},Z$). A nonparametric estimator is preferable, allowing $X$ to take on any values and arbitrarily depend on ($X^{*}$, $Z$).
[40] proposed a semiparametric approach to logistic regression with outcome and exposure errors, which nonparametrically (i.e., without assuming a particular distribution or structure) estimates the exposure error mechanism as described. The current setting is a special case with only exposure error, and a brief outline is provided in the Supplemental Materials. The logiSieve package contains R functions to fit this model [41]. For full technical details on the semiparametric sieve maximum likelihood estimators (SMLEs), see [40].
2.4 Design: Targeted Sampling to Select Informative Patients for Validation
Based on our analysis plan, we considered various sampling designs to select the most informative $100$ patients for validation. When validating computable phenotypes, simple random sampling (SRS) is often used. Under SRS, $100$ patients would be chosen for chart review with equal probability (Figure 3A). This strategy ignores all available information. Some phenotyping algorithms require an SRS validation study. However, for full-likelihood approaches like SMLEs, any information available for all patients can be used to design the validation study. Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
Figure 3: Diagram of partial validation studies with $N$ patients extracted from the electronic health records (EHR) and $n$ patients chosen for validation via chart review. The least informative strategy to select the validation study is A) simple random sampling. More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model. Within strata, the darker shaded areas indicate which patients would be selected.
When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated logOR on ALI), we start at the end. Beginning with the planned logistic regression model of $Y$ given ($X$, $Z$), we work backward to determine which patients will be most informative. For this reason, the “best” designs vary across different estimators, with model-based ones, like SMLEs, offering the most design flexibility. Estimators that rely on the design to handle missingness, like inverse probability weighting (IPW), require nonzero sampling probability for all patients; model-based estimators do not.
We considered stratified and rank-based designs. Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [42, 34, 43]. Ultimately, simulations based on preliminary validation data guided our selection of the targeted design used in the application (see Section 2.6).
2.4.1 Stratified Designs
Given our binary outcome, the case-control (CC) design is a natural option, wherein $50$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 3B). The balanced case-control (BCC) design further stratifies on covariates, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 3C). Continuous covariates must be discretized to form strata. For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average. Then, validate $25$ patients from each of the four ($Y$, $X_{D}^{*}$) strata. Importantly, we can still fit the model using the continuous values $X^{*}$.
The CC and BCC designs are intuitive and offer efficiency gains over SRS. Still, slightly more complicated designs are guaranteed to be theoretically “optimal;” they choose the number of patients from each stratum to achieve the lowest possible variance for the logOR on ALI [44, 45, 46, 47, 37, 48]. [43] proposed the optimal (OPT) design for model-based analysis of data with binary outcome and exposure misclassification (Figure 3D), which could apply here. However, we would achieve the lowest possible variance for $X_{D}^{*}$ rather than $X^{*}$. Increasing the number of unique values of $X_{D}^{*}$ (i.e., more strata) can help, but information will still be lost. Moreover, the OPT design depends on the true logORs, which are unknown; reliable placeholders for them are required to implement the design (often collected in a multi-wave framework).
2.4.2 Rank-Based Designs
Designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains. For example, extreme tail sampling (ETS) designs rank patients by some value of interest and target those with the smallest or largest (i.e., most “extreme”) values. To identify patients most informative not just about ALI but its relationship to healthcare utilization, residual sampling (RS) is preferred [49]. To implement RS with measurement error, we begin by fitting the “naive” analysis model, i.e., using standard logistic regression procedures with $Y$, $X^{*}$ (in place of $X$), and $Z$ to obtain estimates for the logORs $\hat{\boldsymbol{\beta}}^{*}$ from the full sample of $1000$ patients. Then, the naive model residuals are computed for each patient $i$ as
$r_{i}^{*}$ $=Y_{i}-{\Pr}_{\hat{\boldsymbol{\beta}}^{*}}(Y_{i}=1|X^{*}_{i},Z_{i})$
$=Y_{i}-\left[1+\exp\left\{-\left(\hat{\beta}_{0}^{*}+\hat{\beta}_{1}^{*}X_{i}^{*}+\hat{\beta}_{2}^{*}Z_{i}\right)\right\}\right]^{-1},$ (4)
and used to rank patients. The largest residuals $r_{i}^{*}$ (in magnitude) correspond to patients whose observed outcome $Y_{i}$ had a low predicted probability (Figure 3E). The auditDesignR package contains R functions to sample from all of these designs [50].
2.5 Protocol: Enriched Validation Protocol to Reduce Errors and Missingness
Chart reviews were conducted over $6$ months by four clinical research assistants and technicians at WFUSM (hereafter, “auditors”), chosen for their familiarity with Epic (the institution’s electronic charting software) and demonstrated research experience.
2.5.1 Identifying Sources of Auxiliary Information
Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1). While it is unlikely that we would identify the original missing measurements, we harnessed clinical expertise to identify auxiliary information about missing discretized components. The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values. It employs “anchors” to try and recover missing components using clinically-driven placeholders [15].
For example, if a patient were missing HbA1c, auditors would search their chart for a diagnosis of diabetes or impaired glycemic control. If either was present, the patient’s diagnosis would suggest that their missing HbA1c value would have exceeded the $6.5\%$ threshold. Thus, their validated ALI $X$ would include one point for HbA1c. If neither diagnosis was present, $X$ would continue to exclude this still-missing component.
Many components, like HbA1c, had multiple sources of auxiliary information, while others, like triglycerides, had just one. Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
2.5.2 Validated Data Collection
In preparation for validation, the analytical team transformed the extracted EHR data from the original wide format (one row per patient encounter and columns for different variables) to a long format (one row per patient encounter per variable). This transformation streamlined the data structure for auditors: one value to review per row (Supplemental Figure S2).
Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one in Epic). Depending on the scenario, auditors would enter different information for the reviewed value, and each data point was assigned one of five auditor findings (Figure 4).
Figure 4: Flow chart of instructions given to auditors for the chart reviews (Steps 1–3) and the corresponding auditor finding assigned by the analysts (Step 4). Column names refer to those in Supplemental Figure S2.
2.6 Multi-Wave Strategy: Sampling Adaptively and Improving Tools
Choosing the “best” design often relies on information that is unknown at baseline, like EHR error rates or the magnitude of the association between ALI and healthcare utilization. Since we did not have any prior information about these unknowns, we started by sampling conservatively to collect preliminary information for a more extreme, targeted design later on. Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [42, 34]. Even if prior information is available (e.g., from previous chart reviews), multi-wave validation allows us to relearn how data have changed before dedicating resources to validate them [43]. Further, we can seize opportunities to incorporate auditors’ feedback and improve the validation protocol.
We validated $100$ patients’ data across three waves: the Pilot, Wave I, and Wave II. The small Pilot wave tested the validation protocol and established inter-auditor agreement on the first four patients. Then, Waves I and II put this protocol into action, as auditors validated the remaining patients. Waves I and II each selected $48$ more patients. (Designating equal-sized waves balances design flexibility and stability [42].) In Wave II, information from the Pilot + Wave I was leveraged to design a more targeted strategy. Patients could be chosen for validation at most once.
2.6.1 Adapting the Validation Study Design as Information Accumulates
Between the Pilot and Wave I, the first $52$ patients were selected for chart review via the BCC design. Specifically, patients were split into four strata based on their outcome and error-prone ALI (discretized at the median of $0.33$). Then, equal numbers of patients were randomly chosen from the four strata. In the Pilot, one patient was chosen from each stratum; another $12$ were chosen per-stratum in Wave I.
Before validating the remaining $48$ patients, we designed bespoke simulation studies based on the preliminary validation data. They explored how designs compared under different scenarios for true underlying data quality. The design offering the best efficiency (i.e., lowest variability) in estimating the logOR on ALI with the SMLE was desired.
Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency than SRS. The RS design provided the best efficiency under extra low to moderate error settings, and the preliminary validation data (Pilot + Wave I) indicated a perfect true positive rate (TPR) $=1.00$ and negligible false positive rate (FPR) $<0.01$. Under full/high data recovery rates ($90$-$100\%$ missing data recovered through validation) and different designs, the SMLE was empirically unbiased. While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on unvalidated EHR data). Regardless of how much missing data was recovered, the RS design offered the highest efficiency. Details like the data-generating mechanisms can be found in the Supplemental Materials.
We concluded that the RS design was the most informative way to select the remaining $48$ patients for validation in Wave II. To implement it, each unvalidated patient’s residual was calculated based on the naive healthcare utilization model estimates (Section 3.3) following (4), as
$r_{i}^{*}=Y_{i}-\frac{1}{1+\exp\left\{-\left(-1.383+0.945X_{i}^{*}+0.103Z_{i}\right)\right\}}.$
Then, patients were ranked by $r_{i}^{*}$, and the $24$ patients with the smallest residuals and the $24$ patients with the largest residuals were selected for validation. Here, the RS design prioritized (i) patients with zero hospitalizations or ED visits but high predicted probabilities of having at least one and (ii) patients with at least one hospitalization or ED visit but high predicted probabilities of having zero. That is, patients whose observed $Y_{i}$ aligned well with the naive model’s prediction were not considered to be very informative and thus were deprioritized for validation.
2.6.2 Testing the Validation Protocol and Establishing Inter-Auditor Agreement
The validation protocol evolved slightly throughout the waves, as well. After completing the first four patients’ chart reviews (Pilot), auditors and analysts met to discuss any protocol modifications that would be helpful before proceeding. For example, the auditor finding “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded. Adding an optional “Notes” column, where auditors could include any additional information, also resulted from this discussion.
The Pilot also presented a valuable opportunity to evaluate inter-auditor agreement. We started with two auditors on our team. Each reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample. Later, two new auditors joined the team, each reviewing these four patients’ records as well. Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [51], as detailed in the Supplemental Materials. We found high agreement between the auditors’ findings for labs ($\kappa=0.90$) and vitals ($\kappa=0.81$).
3 Results
3.1 Cohort Description and Study Objectives
We considered adult patients ($18$-$65$ years old) with at least one primary care outpatient encounter at Atrium Health Wake Forest Baptist Hospital. The $1000$ sampled patients had a median age of $48$ years and a median income of \$52,320. They were $61\%$ female, $72\%$ White, and $6\%$ Hispanic. During the $2$-year study period, $318$ patients ($32\%$) had at least one hospitalization or ED visit. In unvalidated EHR data, error-prone ALI was right-skewed (Supplemental Figure S3). With a median of $0.33$ (interquartile range $=0.17,0.50$), patients experienced one out of every three non-missing components on their whole-person health, on average. We validated ALI data for $100$ patients.
First, we discuss the EHR data quality (Section 3.2). Second, we present the healthcare utilization model (Section 3.3). Within these sections, results are discussed chronologically, beginning with findings from the unvalidated EHR data and ending with those following the completion of all validation waves. We also discuss how these findings inform future implementations of the ALI in this healthcare system (Section 3.4).
3.2 Validation Study Findings: EHR Data Quality
As expected, the extracted EHR data had lots of missingness. Rare labs for inflammation components were missing for $>95\%$ of patients (Supplemental Figure S4A). Labs for metabolic components were missing for $21\%$ of patients. Common patterns of missing data are displayed in Supplemental Figure S5. Measurements for the cardiovascular components were non-missing for basically everyone. Fortunately, most patients ($78\%$) had at least five non-missing ALI components (Supplemental Figure S4B).
3.2.1 Pilot + Wave I Chart Reviews
[…]
Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined. These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable. Auditor findings were mutually exclusive and specific to whether the measurements were originally non-missing (“extracted value correct,” “extracted value incorrect,” or “extracted value not found”) or missing (“no auxiliary information found” or “auxiliary information found”) in the extracted EHR data.
Figure 6: Comparison of the discretized allostatic load index components from the unvalidated EHR data to A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined.
Fewer vitals than labs matched Epic ($59\%$ versus $81\%$), with $31\%$ not found. We discovered that the EHR data extraction algorithm expected every patient encounter to have vitals. For encounters where vitals were not taken (e.g., physical therapy), values were imputed from other encounters; imputed values could not be found in Epic by the auditors. Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
Incorporating the audit roadmap into the validation protocol improved the completeness of the data. The number of missing values decreased for nearly all ALI components (Supplemental Figure S4A). Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of $52$ validated patients ($44\%$). The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average. Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).
3.2.2 Wave II Chart Reviews
In Wave II, $3867$ data points were validated ($718$ labs, $3149$ vitals), which was fewer than in the Pilot + Wave I. This difference was explained by fewer patient encounters, with $16$-$453$ validated data points per patient (median $=65.5$). The estimated error rates for non-missing components (TPR $=99\%$, FPR $=3\%$) and missing data recovery rate ($17\%$) in Wave II closely resembled the Pilot + Wave I (Figure 6B). Still, some auditors’ findings differed (Figure 5B). For example, more vitals matched Epic ($65\%$).
Combined findings from all validation waves indicated that non-missing EHR data were of excellent quality (TPR $=99.5\%$, FPR $=2\%$), and auxiliary information could be located for $21\%$ of missing components (Figures 5C and 6C). On average, the number of non-missing components per patient increased from $6$ to $7$. Supplemental Tables S1–S2 contain detailed findings.
3.3 Analysis Findings: Whole-Person Health and Healthcare Utilization
Using ALI values from the EHR data, we fit first the naive healthcare utilization model. The expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 were $0.25$ (95% confidence interval [95% CI]: $0.18$, $0.36$). For every $0.1$-point (one-component) increase in ALI, these odds were expected to increase by $10\%$ (OR $=1.10$, 95% CI: $1.03$, $1.18$), adjusting for age. For every $10$ years older a patient was, their odds were expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI. While these findings align with the literature, we wanted to ensure that missingness and potential errors in the EHR data did not lead to spurious conclusions.
3.3.1 Preliminary Estimates Based on the Pilot + Wave I Validation
Incorporating the Pilot + Wave I, we re-fit the healthcare utilization model with the first $52$ patients’ validated data and all $1000$ patients’ unvalidated data. These “preliminary SMLEs” were used to design the simulations in Section 2.6.1. Here, we compare them to the naive model. For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$). The association between whole-person health and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase in ALI (OR $=1.21$, 95% CI: $1.17,1.25$), adjusting for age. The adjusted OR $=1.10$ for age was relatively unchanged (95% CI: $1.01,1.21$). Thus, the preliminary SMLEs reaffirmed that worse whole-person health was associated with higher expected odds of engaging in the healthcare system, perhaps more strongly than the naive model indicated.
3.3.2 Final Estimates Based on All Waves of Validation
After completing all three waves, we used the $100$ patients’ validated data with all $1000$ patients’ unvalidated data to re-fit the model one last time. All final estimates fell between the naive model and preliminary SMLEs. The expected odds of engaging in the healthcare system for an 18-year-old patient with an ALI of 00 were $0.23$ (95% CI: $0.16$, $0.33$). For each $0.1$-point increase in ALI, a $12\%$ increase in odds was expected (OR $=1.12$, 95% CI: $1.09,1.15$), adjusting for age. Again, older patients were expected to be more likely to engage, controlling for ALI (OR $=1.11$, 95% CI: $1.02,1.20$). Additional technical details, including R code, are available in the Supplemental Materials.
3.4 Informing Implementation in the Learning Health System
Ultimately, all models agreed that worse whole-person health (i.e., higher ALI) was associated with a higher likelihood of engaging in the healthcare system, adjusting for age. Chart reviews uncovered very high quality of non-missing EHR data and located auxiliary information for some missing data. Without validation, knowing whether we could trust the EHR data and whether the EHR-derived ALI was still a valid computable phenotype for whole-person health would not have been possible. Thus, while the models led to the same conclusions here, validating EHR data before using them in practice or research is critical. Our findings can now inform implementation strategies to incorporate the ALI into this learning health system.
Having established ALI as a strong predictor of healthcare utilization, embedding it in the EHR is a critical next step. If ALI automatically prepopulated in patients’ electronic charts, clinicians would have easy access to it when making care decisions and researchers could extract it for further studies. This healthcare system previously implemented another score, the EFI, across the patient population aged $65+$. Building on that experience and encouraged by the ALI’s predictive ability, we plan to operationalize it for patients $18$ years and older.
Of course, we want to embed the validated ALI, and only a small subset of patients can undergo chart reviews. Fortunately, in obtaining SMLEs for the healthcare utilization model, we robustly estimated the exposure error mechanism, $\Pr(X|X^{*},Z)$. When scaling up across the healthcare system, the estimated exposure error mechanism can be used to efficiently predict patients’ validated ALI from unvalidated EHR data (Figure 7). See the Supplemental Materials for details on calculating these predictions.
Figure 7: Among patients with chart reviews (triangles), validated allostatic load index (ALI) could differ noticeably from the unvalidated version using the EHR data. For patients without chart reviews (squares), validated ALI could be robustly predicted using the estimated exposure error mechanism. The dashed line indicates the line of equality (i.e., where unvalidated and validated ALI are the same). The solid line is a loess smoother.
4 Discussion
Predicting healthcare utilization is a priority in the learning health system, with important implications at both patient and institutional levels. Establishing standardized, valid measures of whole-person health (like ALI) that are associated with utilization and can scale up to the entire institution is an integral step. While EHR data contain variables to calculate these measures, missing data is a key challenge, and non-missing data are still prone to errors.
Data validation, often through chart reviews, can gauge data quality, but this time-intensive process is usually only possible for a small subset of patients. Also, existing validation protocols do not overcome missingness. We propose novel methods to maximize information gained from a partial validation study of $100$ patients from an EHR dataset of $1000$ through the statistical model, validation protocol, and targeted strategy to select patients. We focused on estimating the association between ALI and healthcare utilization. Partial validation can also efficiently improve estimates of the ALI itself, in preparation for implementation in the EHR.
The proposed analytical and design strategies were devised to ensure robust and efficient estimates. Specifically, the statistical model avoids making restrictive assumptions about the error mechanism, relating the unvalidated and validated ALI, while incorporating all available information to achieve precise estimates (i.e., low variance). Moreover, the targeted design strategies are tailored to the preliminary validation data and chosen statistical model, selecting patients most informative for the clinical question. Our enriched validation protocol builds on the literature (e.g., [52]) by incorporating a clinically-driven “roadmap” to locate auxiliary information about missing values.
The robust SMLEs can handle conditioning on error-prone $X^{*}$ and maybe one other covariate $Z$ in the error mechanism, due to the curse of dimensionality. We assumed that validated ALI depended only on its unvalidated version. However, if a more complicated error structure is expected, data pre-processing will be needed before using the SMLEs Tao2021.
While our roadmap helped recover some missing ALI components, validated patients could still have incomplete data. Clinical guidelines were only defined for missing values presumed to be unhealthy. If auditors were given additional guidance on what auxiliary information might mean that a missing value is healthy, missingness could be further reduced. Others have described missingness in the EHR as selection bias and sought to correct for it using approaches like IPW [53]. Applying such approaches to our post-validation data may reduce the impact of residual missingness. Alternatively, we could adopt other allostatic load calculations, like the index of cardiometabolic health (ICMH), which rely less on rare labs and may face fewer missing components [9].
EHR data are susceptible to “informative presence bias,” wherein the process through which patients engage in the healthcare system can bias associations [54, 55, 56]. Our decision to adapt the Seeman et al. [7] definition of the ALI as the proportion of non-missing components could be considered a summary measure. A recent review found summary measures to be the most common approach to handling informative presence bias [57].
Whole-person health scores like ALI are among many computable phenotypes currently applied to EHR data [16]. Other phenotyping algorithms, including EHRshot [58], also seek to translate observable characteristics into measures of underlying patient states. Validation is a critical step in their development and typically involves chart reviews to obtain some “gold standard” labels [59, 60]. When phenotyping with machine learning methods, this validation set generally needs to be an SRS. Meanwhile, our statistical modeling approach accommodates targeted sampling of the validation study. This flexibility creates opportunities to promote the efficiency of downstream analyses (e.g., associations between the phenotype and suspected risk factors).
There are several interesting directions for future work. First, a longitudinal approach to modeling healthcare utilization, wherein patients’ ALI can vary over time, might provide more information. Second, auditors recorded their findings in a Microsoft Excel workbook. For ease of use and accuracy, an electronic REDCap [61] interface could be built following published guidelines [35]. Third, targeted validation strategies could be used to improve predictive ability, rather than efficiency [62, 63].
5 Conclusion
EHR data provide an exciting opportunity to implement key measures in the learning health system. However, errors and missingness in these data pose challenges in verifying and operationalizing such measures. Validation of even a subset of patients’ data is time- and cost-intensive, but provides crucial insights into data quality. Statistical and biomedical informatics methods can maximize information gained from such partial validation.
Acknowledgments
The authors gratefully acknowledge Wake Forest University and Wake Forest University School of Medicine for an intercampus collaborative grant that supported this work. Computations were performed using the Wake Forest University (WFU) High Performance Computing Facility, a centrally managed computational resource available to WFU researchers including faculty, staff, students, and collaborators.
Supplementary Materials